# Positive Semidefinite Supermartingales and Randomized Matrix Concentration Inequalities

### v3

@@ 4.2 Matrix Generalizations of Classical Test Supermartingales
We delay the proof that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$ to Section B.4. The reason why it is called “empirical” Bernstein is the “plug-in empirical variance” component term $(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}$, as $\widehat{\mathbf{X}}_{n}$ can usually be taken as an empirical average (possibly weighted) from $\mathbf{X}_{1}$ to $\mathbf{X}_{n-1}$, in which case the condition $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$ is satisfied when $\{\mathbf{X}_{n}\}$ all take values in $\mathcal{S}_{d}^{[0,1]}$. We shall come to this empirical Bernstein result again with some scalarization tricks in Section 5.3, where we obtain a strikingly sharp spectral concentration inequality in closed form, further justify its name, discuss the context of empirical Bernstein bounds, and compare it with other existing results.
These supermartingales lead to implicit concentration bounds when applied Theorem 4.4, and are thus unfavorable when one needs a confidence interval or confidence sequence. However, these supermartingales can be used to test whether $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ or not and are thus “test supermartingales” only.
4.3 Matrix e-Processes
Scalar e-processes are an important class of processes generalizing scalar nonnegative supermartingales [Ramdas et al., 2022]. We recall the definition of these processes below, noting that we will always consider a sequence of observations $X_{1},X_{2},\dots$ (while allowing their distribution to vary) and denote their canonical filtration by $\{\mathcal{F}_{n}\}$. We use notations $\mathbb{P}_{P}$, $\mathbb{E}_{P}$ and the general terminology “for $P$” (e.g. we may say a stochastic process adapted to $\{\mathcal{F}_{n}\}$ is a martingale for $P$) to refer to probabilistic statements when the observations $X_{1},X_{2},\dots$ are distributed jointly according to $P$.
@@ 4.3 Matrix e-Processes
Let $\mathcal{P}$ be a set of distributions over $\mathbb{R}\times\mathbb{R}\times\dots$. A nonnegative stochastic process $\{Y_{n}\}$ adapted to $\{\mathcal{F}_{n}\}$ is called an e-process for $\mathcal{P}$ if, for any $P\in\mathcal{P}$, there exists a supermartingale $\{M_{n}^{P}\}$ adapted to $\{\mathcal{F}_{n}\}$ with $M_{0}^{P}\leqslant 1$ that upper bounds the process $\{Y_{n}\}$. Equivalently, $\{Y_{n}\}$ is an e-process if $\mathbb{E}_{P}[Y_{\tau}]\leqslant 1$ for any stopping time $\tau$ on $\{\mathcal{F}_{n}\}$ [Ramdas et al., 2020, Lemma 6]. Clearly, if $\{Y_{n}\}$ is a nonnegative supermartingale for $\mathcal{P}$ (meaning for every $P\in\mathcal{P}$), the optional stopping theorem implies that it is also an e-process for $\mathcal{P}$, but e-processes are a much more general class [Ramdas et al., 2022, Ruf et al., 2023].
A scalar e-process satisfies the following form of (VI) or (UVI):
for $\text{for any }P\in\mathcal{P},\mathbb{P}_{P}(Y_{\tau}\geqslant Ua)\leqslant a^{-1}$
@@
for any stopping time $\tau$ and independent super-uniform random variable $U$ (and thus we also have that $\mathbb{P}_{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}$ by setting $U=1$). This property enables scalar e-processes to be used for testing a hypothesis that the data is drawn from some $P\in\mathcal{P}$: we may reject the null as soon as the e-process exceeds $1/\alpha$ (for some prespecified constant $\alpha\in(0,1)$) and be guaranteed that the type-I error is at most $\alpha$. E-processes are particularly useful for composite sequential testing problems for which nonnegative supermartingales are powerless [Ramdas et al., 2022, Ruf et al., 2023].
We can generalize the scalar e-processes to our matrix setting as well, with the sequence of observations that generates the filtration $\{\mathcal{F}_{n}\}$ now denoted by $\mathbf{X}_{1},\mathbf{X}_{2},\dots$.
@@
Let $\mathfrak{P}$ be a set of distributions over $\mathcal{S}_{d}\times\mathcal{S}_{d}\times\dots$. An $\mathcal{S}_{d}^{+}$-valued stochastic process $\{\mathbf{Y}_{n}\}$ adapted to $\{\mathcal{F}_{n}\}$ is called a matrix e-process for $\mathfrak{P}$ if, for any $\mathfrak{p}\in\mathfrak{P}$, there exists an $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{M}_{n}^{\mathfrak{p}}\}$ adapted to $\{\mathcal{F}_{n}\}$ with $\mathbf{M}_{0}^{\mathfrak{p}}\preceq\mathbf{I}$ that upper bounds the process $\{\mathbf{Y}_{n}\}$, i.e., $\mathbf{M}_{n}^{\mathfrak{p}}\succeq\mathbf{Y}_{n}$ for all $n$ almost surely for $\mathfrak{p}$.
Similar to the scalar case, matrix e-processes satisfy their (UMVI).
Lemma 4.12 (Uniformly Randomized Matrix Ville’s Inequality for Matrix e-Processes).
Let $\{\mathbf{Y}_{n}\}$ be a matrix e-process for $\mathfrak{P}$. Then, for any $\mathfrak{p}\in\mathfrak{P}$, for any stopping time $\tau$ on $\{\mathcal{F}_{n}\}$, any trace super-uniform random matrix $\mathbf{U}$ taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ and any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
@@
$\mathbb{P}_{\mathfrak{p}}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1}).$ (40)
Proof.
Consider the upper bounding matrix supermartingale $\{\mathbf{M}_{n}^{\mathfrak{p}}\}$. By (UMVI), $\mathbb{P}(\mathbf{M}_{\tau}^{\mathfrak{p}}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1})$. Since $\mathbf{M}_{n}^{\mathfrak{p}}\succeq\mathbf{Y}_{n}$ almost surely for any $n$, $\mathbf{M}_{\tau}^{\mathfrak{p}}\succeq\mathbf{Y}_{\tau}$ almost surely as well. Therefore, $\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathbb{P}(\mathbf{M}_{\tau}^{\mathfrak{p}}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1}).$
@@
Simple examples of matrix e-processes following the previous subsection include lower bounds of $\mathcal{S}_{d}^{+}$-valued supermartingales. First, we define the following “matrix minimum”: for $\mathbf{A},\mathbf{B}\in\mathcal{S}_{d}$, we define $\mathbf{A}\curlyvee\mathbf{B}$ to be $\mathbf{A}$ if $\mathbf{A}\preceq\mathbf{B}$, and $\mathbf{A}-\lambda_{\max}(\mathbf{A}-\mathbf{B})\cdot\mathbf{I}$ if otherwise. Note that $\mathbf{A}\curlyvee\mathbf{B}\preceq\mathbf{A}$ and $\mathbf{A}\curlyvee\mathbf{B}\preceq\mathbf{B}$, and we remark that in general there is no largest symmetric matrix lower bound of $\mathbf{A}$ and $\mathbf{B}$ in the $\preceq$ order [Kadison, 1951, Theorem 6].
Let $\mathbf{M},\mathbf{M}^{\prime}\preceq\mathbf{B}$ be three matrices in $\mathcal{S}_{d}^{++}$, and we define $\mathcal{B}(\mathbf{B},\mathbf{M})$ to be the set of distributions on $\mathcal{S}_{d}^{+}$ upper bounded by $\mathbf{B}$ with mean $\mathbf{M}$. Further, we define the distribution families
$\mathfrak{P}=\{P_{1}\otimes P_{2}\otimes\dots:P_{n}\in\mathcal{B}(\mathbf{B},\mathbf{M})\text{ for all }n\}$ (41)
@@
$\mathfrak{P}^{\prime}=\{P_{1}\otimes P_{2}\otimes\dots:P_{n}\in\mathcal{B}(\mathbf{B},\mathbf{M}^{\prime})\text{ for all }n\}.$ (42)
According to Example 4.7, we can construct a $\mathcal{S}_{d}^{+}$-valued betting process $\{\mathbf{Y}_{n}\}$ which is a supermartingale under any of $\mathfrak{p}\in\mathfrak{P}$, and another $\{\mathbf{Y}_{n}^{\prime}\}$ which is a supermartingale under any of $\mathfrak{p}\in\mathfrak{P}^{\prime}$. Then the minimum process $\{\mathbf{Y}_{n}\curlyvee\mathbf{Y}_{n}^{\prime}\}$ is a matrix $\mathfrak{P}\cup\mathfrak{P}^{\prime}$-e-process, useful for testing sequentially the null that the true mean is always either $\mathbf{M}$ or $\mathbf{M}^{\prime}$.
4.4 Matrix-Valued Submartingales and Backward Supermartingales
Next, we note that all the above are still valid if we replace “supermartingale”s with “submartingale”s. To wit, we say that $\{\mathbf{Y}_{n}\}$ is a submartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq\mathbf{Y}_{n-1}$. It is clear from the proof of Lemma 4.3 that,

### v4

@@ 3.3 Examples: Conditional Mean Testing Problems
Let us briefly explain the name “empirical Bernstein”. The original scalar version of the inequality above (that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$) was proved by Howard et al. [2021, Theorem 4], generalized and studied closer by Waudby-Smith and Ramdas [2023, Theorem 2], producing a confidence sequence for the common conditional mean of martingale dependent, $[0,1]$-valued scalars that is asymptotically as sharp as the Bernstein inequality without assuming knowing the underlying variance while adapting to it, and beats numerous then-existing empirical Bernstein inequalities in the literature. The matrix inequality that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$ above was recently proved by Wang and Ramdas [2024, Lemma 4.1], who applied the same scalarization technique as Howard et al. [2020] (see Section 5.2) to obtain a scalar test supermartingale and a closed-form spectral empirical Bernstein inequality for $\mathcal{S}_{d}^{[0,1]}$-valued random matrices, of similar sharpness.
Here, on the other hand, we demonstrate that if a closed-form spectral confidence set is not of interest, one can sequentially test the null $\mathbb{E}[\mathbf{X}_{n}|\mathcal{F}_{n-1}]=\mathbf{M}$, under the assumption that $\{\mathbf{X}_{n}\}$ all take values in $\mathcal{S}_{d}^{[0,1]}$, by constructing a matrix-valued test supermartingale using Lemma 3.5 of potential larger power compared to the scalarization approach of Wang and Ramdas [2024]. The hyperparameters (predictable scalars $\{\gamma_{n}\}$ and matrices $\{\widehat{\mathbf{X}}_{n}\}$) in Example 3.10 can be set according to the discussion of Wang and Ramdas [2024, Section 4].
Our last example generalizes the scalar “Catoni supermartingales” due to Catoni [2012] in the fixed-time case and Wang and Ramdas [2023a] in the sequential case. With scalars, this influence function-based technique produces subGaussian confidence intervals and sequences for heavy-tailed random variables with a $p$th moment upper bound where $p\in(1,2]$. In the $p=2$ special case Wang and Ramdas [2023a] demonstrate that the “Catoni-style confidence sequence” is tighter than the self-normalization technique which we already generalized to matrices in Example 3.8, under a variance upper bound assumption. Our matrix extension of these results starts with the standard definition of Catoni’s influence function (see e.g. Chen et al. [2021, Section 1], Wang and Ramdas [2023a, Section 9]).
Definition 3.11.
An increasing function $\phi:\mathbb{R}\to\mathbb{R}$ is upper $p$-logarithmically contractive if $\phi(x)\leqslant\log(1+x+|x|^{p}/p)$; lower $p$-logarithmically contractive if $-\log(1-x+|x|^{p}/p)\leqslant\phi(x)$.
For example, the function
$\phi_{p}(x)=\begin{cases}\log(1+x+x^{p}/p),&x\geqslant 0;\\ -\log(1-x+(-x)^{p}/p),&x<0.\end{cases}$ (27)
is both upper and lower $p$-logarithmically contraction.
Example 3.12 (Matrix Catoni Supermartingales).
Let $\{\mathbf{X}_{n}\}$ be an adapted sequence of $\mathcal{S}_{d}$-valued random matrices with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional $p$th moment upper bounds $\mathbb{E}((\operatorname{abs}(\mathbf{X}_{n}-\mathbf{M}_{n}))^{p}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{n}$ that are predictable. For any upper $p$-logarithmically contractive function, we can take in Lemma 3.5,
$\mathbf{E}_{n}=\exp(\phi(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n}))),\quad\mathbf{F}_{n}=\left(\mathbf{I}+\frac{\gamma_{n}^{p}}{p}\mathbf{V}_{n}\right)^{-1};$ (28)
and for any lower $p$-logarithmically contractive function, we can take in Lemma 3.5,
$\mathbf{E}_{n}=\exp(-\phi(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n}))),\quad\mathbf{F}_{n}=\left(\mathbf{I}+\frac{\gamma_{n}^{p}}{p}\mathbf{V}_{n}\right)^{-1}.$ (29)
Above, $\{\gamma_{n}\}$ are predictable positive scalars.
Example 3.12 is proved in Section B.2. We remark that, in the scalar case, the two supermartingales constructed this way remain supermartingales under one-sided nulls, one for each side. This fails to hold with matrices (e.g., the null that $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{M}_{n}$) as there is no guarantee that $\exp(\pm\phi(\cdot))$ are operator monotone. Still, the two matrix test supermartingales play the role of detecting null-deviating behavior of opposite directions, and a powerful test for the point null $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ can be formed by a convex combination of the two.
3.4 Matrix e-Values and e-Processes
We have thus far generalized scalar nonnegative supermartingales to PSD matrix supermartingales, as well as their numerous applications in sequential statistics as “test supermartingales”. However, with scalars, it is known that nonnegative supermartingales fail to be powerful in some examples involving large composite null hypotheses, and a wider class of processes called e-processes is often used instead [Ramdas et al., 2022]. Loosely speaking, an e-process is a nonnegative process upper bounded by supermartingales. We first recall the formal definitions of scalar e-values and e-processes, and generalize them to matrices as well.
Throughout this subsection, we consider $X_{1},X_{2},\dots$ a sequence of random observations taking values in some space $\mathcal{X}$ and they generate the filtration $\{\mathcal{F}_{n}\}$. We use notations $\mathbb{P}_{P}$, $\mathbb{E}_{P}$ and the general terminology “for $P$” (e.g. we may say an adapted stochastic process is a martingale for $P$) to refer to probabilistic statements when the observations $X_{1},X_{2},\dots$ are distributed jointly according to $P$, where $P$ is some distribution on $\mathcal{X}\times\mathcal{X}\times\dots$.
@@ 3.4 Matrix e-Values and e-Processes
Let $\mathcal{P}$ be a set of distributions on $\mathcal{X}\times\mathcal{X}\times\dots$. An $\mathcal{F}_{\infty}$-measurable nonnegative random variable $Y$ is called an e-value for $\mathcal{P}$ if, for any $P\in\mathcal{P}$, its mean $\mathbb{E}_{P}Y\leqslant 1$. An adapted nonnegative process $\{Y_{n}\}$ is called an e-process for $\mathcal{P}$ if, for any stopping time $\tau$, the stopped variable $Y_{\tau}$ is an e-value for $\mathcal{P}$. It is known that, due to Ramdas et al. [2020, Lemma 6], $\{Y_{n}\}$ is an e-process if and only if for any $P\in\mathcal{P}$ it is upper bounded by a supermartingale $\{M_{n}^{P}\}$ with $M_{0}^{P}\leqslant 1$. Clearly, if $\{Y_{n}\}$ is a nonnegative supermartingale for $\mathcal{P}$ (i.e. for every $P\in\mathcal{P}$), the optional stopping theorem implies that it is also an e-process for $\mathcal{P}$, but e-processes are a much more general class [Ramdas et al., 2022, Ruf et al., 2023].
Scalar e-processes also satisfy (VI), both time-uniformly and at a stop time:
$\text{For any $P\in\mathcal{P}$}:\quad\mathbb{P}_{P}(\exists n,\ Y_{n}\geqslant a)\leqslant a^{-1};$ (30)
and for any stopping time $\text{and for any stopping time $\tau$},\quad\mathbb{P}_{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}.$ (31)
@@
This property enables scalar e-processes to also be used for testing the null hypothesis that the data is drawn from some $P\in\mathcal{P}$: we may reject the null as soon as the e-process exceeds $1/\alpha$ (for some prespecified constant $\alpha\in(0,1)$) and be guaranteed that the type-I error is at most $\alpha$. E-processes are particularly useful for composite sequential testing problems for which nonnegative supermartingales are powerless [Ramdas et al., 2022, Ruf et al., 2023].
We define matrix e-values and e-processes similarly.
Definition 3.13 (Matrix e-Value).
Let $\mathcal{P}$ be a set of distributions on $\mathcal{X}\times\mathcal{X}\times\dots$. An $\mathcal{S}_{d}^{+}$-valued, $\mathcal{F}_{\infty}$-measurable random matrix $\mathbf{Y}$ is called a (matrix) e-value for $\mathcal{P}$ if $\mathbb{E}_{P}\mathbf{Y}\preceq\mathbf{I}$ for any $P\in\mathcal{P}$.
@@
Let $\mathcal{P}$ be a set of distributions on $\mathcal{X}\times\mathcal{X}\times\dots$. An $\mathcal{S}_{d}^{+}$-valued adapted stochastic process $\{\mathbf{Y}_{n}\}$ is called a (matrix) e-process for $\mathcal{P}$ if, for any stopping time $\tau$, the stopped random matrix $\mathbf{Y}_{\tau}$ is a matrix e-value for $\mathcal{P}$.
It can be seen from the matrix optional stopping theorem (Theorem 3.3) that if $\{\mathbf{Y}_{n}\}$ is a PSD supermartingale for $\mathcal{P}$ then it is also an e-process for $\mathcal{P}$. More generally, $\{\mathbf{Y}_{n}\}$ is a matrix e-process for $\mathcal{P}$ if, for any $P\in\mathcal{P}$, there exists an $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{M}_{n}^{P}\}$ with $\mathbf{M}_{0}^{P}\preceq\mathbf{I}$ that upper bounds the process $\{\mathbf{Y}_{n}\}$, i.e., $\mathbf{M}_{n}^{P}\succeq\mathbf{Y}_{n}$ for all $n$ almost surely for $P$. Similar to the scalar case above, matrix e-processes satisfy (MVI), both time-uniform and at a stopping time:
@@
$\text{For any $P\in\mathcal{P}$}:\quad\mathbb{P}_{P}(\exists n,\ \mathbf{Y}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{A}^{-1});$ (32)
and for any stopping time $\text{and for any stopping time $\tau$},\quad\mathbb{P}_{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{A}^{-1}),$ (33)
which follows directly from the proof of Theorem 3.4. Therefore, one can run a procedure similar to Algorithm 1 with an e-process instead for sequential hypothesis testing.
@@
Simple examples of matrix e-processes following the previous subsection therefore include lower bounds of $\mathcal{S}_{d}^{+}$-valued supermartingales. In the scalar case, an e-process is often constructed for a composite null hypothesis by taking an infimum over the test supermartingales under each of the null distributions. While a similar construction carries over to the matrices, it is important to note that in general the matrix minimum does not exist: for most pairs $\mathbf{A},\mathbf{B}\in\mathcal{S}_{d}$, there is no largest lower bound of $\mathbf{A}$ and $\mathbf{B}$ in the $\preceq$ order within $\mathcal{S}_{d}$ [Kadison, 1951, Theorem 6]. However, we can define a matrix minimum, for example as follows:
Definition 3.15.
We define the following minimum operation $\curlyvee:\mathcal{S}_{d}\times\mathcal{S}_{d}\to\mathcal{S}_{d}$:
$\mathbf{A}\curlyvee\mathbf{B}=\begin{cases}\mathbf{A},&\text{if }\mathbf{A}\preceq\mathbf{B};\\ \mathbf{A}-\lambda_{\max}(\mathbf{A}-\mathbf{B})\cdot\mathbf{I},&\text{otherwise}.\end{cases}$ (34)
Note that $\mathbf{A}\curlyvee\mathbf{B}\preceq\mathbf{A}$ and $\mathbf{A}\curlyvee\mathbf{B}\preceq\mathbf{B}$, and that in general $\mathbf{A}\curlyvee\mathbf{B}\neq\mathbf{B}\curlyvee\mathbf{A}$.
[…]
$\mathbf{C}\preceq\mathbf{A},\mathbf{C}\preceq\mathbf{B},\mathbf{C}\npreceq\mathbf{A}\curlyvee\mathbf{B}.$ (35)
For example, $\mathbf{B}\curlyvee\mathbf{A}$ is often an example of $\mathbf{C}$ that satisfies above. On the other hand, however, it is easy to verify that there does not exist a matrix $\mathbf{C}$ that satisfies
$\mathbf{C}\preceq\mathbf{A},\mathbf{C}\preceq\mathbf{B},\mathbf{C}\succ\mathbf{A}\curlyvee\mathbf{B},$ (36)
in the sense that there are several choices of “matrix minima”, Definition 3.15 being one among them, and these choices do not dominate one another. Finally, we note that the matrix minimum enables us to construct matrix e-processes for null hypotheses that are finite unions of several conditions under which a matrix test supermartingale is constructible:
Proposition 3.16.
Suppose $\mathcal{P}=\bigcup_{k=1}^{K}\mathcal{P}_{k}$ and $\{\mathbf{Y}^{k}_{n}\}$ is a matrix e-process for $\mathcal{P}_{k}$. Then, $\{\mathbf{Y}_{n}^{1}\curlyvee\mathbf{Y}_{n}^{2}\curlyvee\dots\curlyvee\mathbf{Y}_{n}^{K}\}$ is a matrix e-process for $\mathcal{P}$.
Here, the $K$-fold minimum $\mathbf{Y}_{n}^{1}\curlyvee\mathbf{Y}_{n}^{2}\curlyvee\dots\curlyvee\mathbf{Y}_{n}^{K}$ can be defined either left-associatively as $(\dots(\mathbf{Y}_{n}^{1}\curlyvee\mathbf{Y}_{n}^{2})\curlyvee\dots)\curlyvee\mathbf{Y}_{n}^{K}$ or right-associatively as $\mathbf{Y}_{n}^{1}\curlyvee(\dots\curlyvee(\mathbf{Y}_{n}^{K-1}\curlyvee\mathbf{Y}_{n}^{K})\dots)$. These lead to two different e-processes. One can take the average over all $K!$ permutations if one desires an e-process that is symmetric w.r.t. all $K$ subclasses. A simple example is as follows. Let $\mathbf{M},\mathbf{M}^{\prime}\preceq\mathbf{B}$ be three matrices in $\mathcal{S}_{d}^{++}$, and we define $\mathcal{B}(\mathbf{B},\mathbf{M})$ to be the set of distributions on $\mathcal{S}_{d}^{+}$ upper bounded by $\mathbf{B}$ with mean $\mathbf{M}$. Further, we define the distribution families
$\mathcal{P}_{1}=\{P_{1}\otimes P_{2}\otimes\dots:P_{n}\in\mathcal{B}(\mathbf{B},\mathbf{M})\text{ for all }n\}$ (37)
@@
$\mathcal{P}_{2}=\{P_{1}\otimes P_{2}\otimes\dots:P_{n}\in\mathcal{B}(\mathbf{B},\mathbf{M}^{\prime})\text{ for all }n\}.$ (38)
According to Example 3.7, we can construct a $\mathcal{S}_{d}^{+}$-valued betting process $\{\mathbf{Y}_{n}\}$ which is a supermartingale under any of $P\in\mathcal{P}_{1}$, and another $\{\mathbf{Y}_{n}^{\prime}\}$ which is a supermartingale under any of $P\in\mathcal{P}_{2}$. Then the minimum process $\{\mathbf{Y}_{n}\curlyvee\mathbf{Y}_{n}^{\prime}\}$ is a matrix e-process for $\mathcal{P}_{1}\cup\mathcal{P}_{2}$, useful for testing sequentially the null that the true mean is always either $\mathbf{M}$ or $\mathbf{M}^{\prime}$.
4 Backward PSD Submartingales
We have so far discussed the properties and applications $\mathcal{S}_{d}^{+}$-valued supermartingales on a forward filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$. We now move on to a backward filtration $\{\mathcal{G}_{0}\}_{n\geqslant 0}$ and its adapted $\mathcal{S}_{d}^{+}$-valued submartingales to extend the established theory with scalars. This involves first considering a forward submartingale of finite length on the finite forward filtration $\{\mathcal{F}_{n}\}_{0\leqslant n\leqslant N}$, inverting the time, and extending indefinitely into the “past” (before the inversion). For this reason, we shall always be explicit with the filtration and the time range in this section.
4.1 Submartingales and Maximal Inequalities: From Forward to Backward
In the proof of Theorem 3.3, we appeal to the simple fact that if $\{\mathbf{Y}_{n}\}$ is a matrix supermartingale then $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a scalar supermartingale, which enables us to derive stopping time theorems for matrix supermartingales via those for scalar supermartingales. This clearly remains still valid if we replace “supermartingale”s with “submartingale”s. To wit, we say that $\{\mathbf{Y}_{n}\}_{n\geqslant 0}$ adapted to $\{\mathcal{F}_{n}\}_{n\geqslant 0}$ is a matrix submartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq\mathbf{Y}_{n-1}$. It is clear from the proof of Theorem 3.3 that,